Sean mows his half of hte yard in 30 minutes. His brother takes 50 minutes. If they each had their own lawn mower and worked together, how long would the job take?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Sean mows his half of the yard in 30 minutes. His brother takes 50 minutes. If they each had their own lawn mower and worked together, how long would the job take?
-------------------------
Sean DATA:
time = 30 min/job ; rate = 1/30 job/min
--------------------------------------------
Brother DATA:
time = 50 min/job ; rate = 1/50 job/min
-------------------------
Together DATA:
time = x min/job ; rate = 1/x job/min
------------------------------
Equation:
rate + rate = together rate
1/30 + 1/50 = 1/x
To get rid of the denominators,
multiply thru by 150x to get:
5x + 3x = 150
8x = 150
x = 150/8 = 18 3/4 minutes (time to do the job together)
=============================================================
Cheers,
Stan H.

Answer by solver91311(24713)



If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period.

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.

Just substitute your given time values and do the arithmetic.


John